I'm having problems putting this into a formula because I feel that something is missing. Please help.
Lance bicycles 37 km/h with no wind. Against the wind he bikes 75 km in the same time it takes to bike 110 km with the wind. What is the speed of the wind?

Answer by josmiceli(19441)

time is the same both with and against the wind
t = time
t = distance/rate
w = wind speed
against the wind his rate is 37 - w
with the wind his rate is 37 + w
t = 75 / (37 - w)
t = 110 / (37 + w)
75 / (37 - w) = 110 / (37 + w)
110*37 - 110*w = 37*75 + 75*w
185*w = 37(110 - 75)
185w = 37*35
w = 37*35 / 185
w = 1295/185
w = 7 km/hr wind speed